OpenStudy (ally_b):

Express the complex number in trigonometric form. -2i

OpenStudy (dumbcow):

\[a +bi = r(\cos \theta + i \sin \theta)\] \[r = \sqrt{a^2+b^2}\] For this example , a = 0 b = -2 ---> r = 2 ---> cos theta = 0 , sin theta = -1 ---> theta = 3pi/2